OpenStudy (anonymous):

local linear approximation to estimate 28^1/3 ?

OpenStudy (triciaal):

one approach 28^1/3 27^1/3 = 3 28^1/3 can be written as (27 + 1 )^1/3 (a + b) ^ n =

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{27+x}\] for \(x=1\)

OpenStudy (anonymous):

use \(f'(0)\) to get local linearization

OpenStudy (anonymous):

at the point \((0,3)\)

OpenStudy (anonymous):

Okay so (1/3)(27+x)^-2/3

OpenStudy (anonymous):

yeah radical form is what is needed here

OpenStudy (anonymous):

so you can evaluate at \(x=0\) to get your slope